OpenStudy (anonymous):

I'm really having a trouble answering this question. Could you please help me.

OpenStudy (anonymous):

Mr. Chan had to arrive at the airport at 4:00 p.m. The airport is 65 km from his house. If he starts at 2:45 p.m. from his house, at what speed should he drive so as to reach the airport on time?

OpenStudy (anonymous):

x=v.t and t is the difference in time so 4.00-2.45= 1h15min=1.25 hours and 65 km so 65=v.1.25 and v=52 km/h

OpenStudy (anonymous):

1h15min=1.25 hours ?

OpenStudy (anonymous):

yes 1h15m = 60+15=75 minutes if you divide 75 by 60, you get 1.25 hours

OpenStudy (anonymous):

O thanks that's the part where i got stuck at.

OpenStudy (anonymous):

yes, that's often the tricky part. First you should notice the units; the problem gives the distance in kilometers so you should solve it with respect to km/h and convert all units of time to hours. Second you should convert minutes to hours.